OpenStudy (anonymous):

I need help simplifying the expression using logarithm properties ln(1/e^5)

OpenStudy (danjs):

hello

OpenStudy (danjs):

recall: ln(a/b) = ln(a) - ln(b)

OpenStudy (danjs):

recall ln(e^x) = x

OpenStudy (anonymous):

so then \[\ln 1-\ln e^5\] ??

OpenStudy (danjs):

yes

OpenStudy (danjs):

= ln(1) - 5

OpenStudy (anonymous):

Thank you!